OpenStudy (anonymous):

which of the following is the minimum value of the equation y=-x^2=2x+5 HELP PLEASEEEEEEEEEEEE!!!

OpenStudy (anonymous):

I supose:y=-x^2-2x+5 y' =-2x-2 make y' = 0 -2x-2=0 x =-1/2 put it in the formula: y=-1/4 +1+5 = 6-1/4 =23/4

OpenStudy (anonymous):

answer key is either 5,6,1 or 2